# Is Limited Information Enough? An Approximate Multi-agent Coverage Control in Non-Convex Discrete Environments

### 08 Jan 2024 v1

@@ 3.1. High Level Description of the Algoritm
Note that determining the locations $y_{1},\dots y_{k}$ maximizing Equation (13) is actually an NP-hard problem, since it consists in general in solving a multiagent coverage problem.
Given this, we can find the agent $i_{\max}^{+}$ which would contribute to $G(x)$ the most from having an additional agent in her region:
$\left.\begin{array}[]{l}i_{\max}^{+}(x,\mathcal{P})=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\mathop{\rm argmax\leavevmode\nobreak\ }\limits}\leavevmode\nobreak\ M_{1}(x_{i};\mathcal{P}_{i}).\end{array}\right.$ (14)
$\left.\begin{array}[]{l}V(x,\mathcal{P})=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\max}\leavevmode\nobreak\ M_{1}(x_{i};\mathcal{P}_{i}),\end{array}\right.$ (15)
@@
(Ex. 1, cont.). We have $i_{min}=a$, since $u_{a}=1$. Furthermore, the agent which would contribute to $G(x)$ the most from adding a single agent within her region is agent $e$, as adding an agent (depicted as ’+’) would induce $G$ of $6.5$ in her region, thus $M_{1}(x_{e},\mathcal{P}_{e})\simeq 6.5-5=1.5$ (See Figure 3 for the illustration).
@@
Figure 3. Agent $e$ would contribute to $G(x)$ the most from the addition of an agent (depicted as ’+’) in her region
Before going into the details, we show the following main result.
Theorem 5 (Convergence with Performance guarantee).
Assuming that agents have access to $u_{\min}$ and $i_{\max}^{+}$ given any configuration, Algorithm 2 converges to a neighborhood optimum and terminates. The approximation ratio of the algorithm is 2.
Note that the approximation ratio of 2 is equal to the lower bound predicted by Equation (8) in case of full information ($f=n$). Now we explain how we construct Algorithm 2 and reach this conclusion. We classify the solution $(x,\mathcal{P})$ into 4 states. First, the solution space is divided into the following two states $Z_{1}$ and $Z_{2}$, by checking if $i_{\max}^{+}$ would gain more from adding an agent in her region:
@@
In $Z_{2}$, no single agent would contribute enough from accommodating $i_{\min}$. Next the solutions in $Z_{2}$ are further classified depending on whether integrating a further agent in the neighborhood could induce a significant enough marginal gain. We denote $n_{i}=|\mathcal{N}_{i}|$ and $\mathcal{P}_{n_{i}}=\cup_{j\in\mathcal{N}_{i}}\mathcal{P}_{j}$. In the following state $Z_{3}$, there is no such a group:
@@
$M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\leq u_{\min}(x,\mathcal{P}),$ (18)
$\leavevmode\nobreak\ \forall i\in\mathcal{N}\},$
Intuitively, $M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})$ is the efficiency of a neighborhood optimum of $\mathcal{N}_{i}$. Equation (18) means that the gain in the neighborhood optimum by accommodating $i_{\min}$ in $\mathcal{P}_{n_{i}}$ cannot be larger than $u_{\min}$.
@@
$\underset{\begin{subarray}{c}j\in\mathcal{N}_{i}\end{subarray}}{\sum}u_{j}(x,\mathcal{P})=M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}}),\leavevmode\nobreak\ \forall i\in\mathcal{N}\}.$ (20)
@@
(Ex. 1, cont.) The allocation $x$ depicted is in state $Z_{1}$ since $V(x,\mathcal{P})=1.5>1=u_{a}$, with $i_{min}=a$.
Algorithm 1 Communication tree
1:procedure $D_{i}=$CommTree($i$)
2:     $u_{{\rm lmin}}=\hat{u}_{i}$, $j_{{\rm lmin}}=i$
3:     for $j\in\mathcal{\hat{N}}_{i}$  do
4:         if $\hat{u}_{j}<u_{{\rm lmin}}$ then
5:              $u_{{\rm lmin}}=\hat{u}_{j}$, $j_{{\rm lmin}}=j$
6:     $f_{updated}=False$
7:     if $j_{{\rm lmin}}=i$  then
8:         if $D_{i}<0$ then
9:              $D_{i}=i$, $f_{updated}=True$
10:     else
11:         if $D_{i}\not=j_{{\rm lmin}}$ then
12:              $D_{i}=j_{{\rm lmin}}$
13:              $\hat{u}_{i}=u_{{\rm lmin}}$, $f_{updated}=True$
14:     if $f_{updated}$ then
15:         for $j\in\mathcal{\hat{N}}_{i}$  do
16:              $D_{j}=\textsc{CommTree}(j)$
17:     Return $D_{i}$
@@
[…]
Figure 4. (Top left): the neighborhood graph in Example 1. The numbers represent utilities. (Top right): agent ’c’ points to the agent ’b’ that has minimum utility among neighborhood. It also updates its utility. (Bottom left): then agents ’b’ and ’e’ point to the minimum neighbors. (Bottom right): the communication tree in convergence. In this case, agent ’a’ is $i_{\min}$ and therefore the root of the tree.
@@
4:         Communicate $u_{\min}(x,\mathcal{P})$, $x_{i_{\min}}(x,\mathcal{P})$ and $i_{\max}^{+}(x,\mathcal{P})$
5:         Pick $i\in\mathcal{N}$
6:         $\mathcal{N}^{2}_{i}=\textsc{Neighbor}(\mathcal{P},i)$
7:         while $(x,\mathcal{P})\in Z_{1}$ and $i_{\max}^{+}\not\in\mathcal{N}^{2}_{i}$  do
8:              pick $i\in\mathcal{N}$
9:         if $i_{\min}(x,\mathcal{P})\in\mathcal{N}^{2}_{i}$ or
10:            $M_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})-M_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})\leq u_{\min}(x,\mathcal{P})$  then
11:               $x\leftarrow B_{n_{i}}(\emptyset,\mathcal{P}_{n_{i}})$ (Step a)
12:              $\mathcal{P}_{j}\leftarrow\mathcal{V}_{j}(x^{\prime},\mathcal{P}_{n_{i}}),\leavevmode\nobreak\ \forall j\in\mathcal{N}^{2}_{i}$
13:         else
14:              Consider a virtual agent $l\not\in\mathcal{N}$ (Step b)
15:              $x^{\prime}\leftarrow B_{n_{i}+1}(\emptyset,\mathcal{P}_{n_{i}})$
16:              $\mathcal{P}^{\prime}_{k}\leftarrow\mathcal{V}_{k}(x^{\prime},\mathcal{P}_{n_{i}}),\leavevmode\nobreak\ \forall k\in\mathcal{N}^{\prime}_{i}=\{1,\ldots,n_{i}+1\}$
@@
17:              $\tilde{i}_{\min}=\underset{\begin{subarray}{c}j\in\mathcal{N}_{\mathcal{N}^{2}_{i}}\end{subarray}}{\mathop{\rm argmin\leavevmode\nobreak\ }\limits}|x_{j}-x_{i_{\min}}|$
@@
18:              $k_{l}=\underset{\begin{subarray}{c}k\in\mathcal{N}^{\prime}_{i}|{\rm adj}(\mathcal{P}^{\prime}_{k},\mathcal{P}_{\tilde{i}_{\min}})\end{subarray}}{\mathop{\rm argmin\leavevmode\nobreak\ }\limits}|x^{\prime}_{k}-x_{\tilde{i}_{\min}}|,\leavevmode\nobreak\ x_{l}\leftarrow x^{\prime}_{k_{l}}$
19:              $x\leftarrow x^{\prime}\setminus\{x_{l}\}$
@@
20:              $i_{+}\leftarrow\underset{\begin{subarray}{c}j\in\mathcal{N}^{2}_{i}|{\rm adj}(\mathcal{P}^{\prime}_{j},\mathcal{P}^{\prime}_{l})\end{subarray}}{\mathop{\rm argmin\leavevmode\nobreak\ }\limits}\leavevmode\nobreak\ |x_{j}-x_{l}|,\leavevmode\nobreak\ \mathcal{P}_{i_{+}}\leftarrow\mathcal{P}_{i_{+}}\cup\mathcal{P}_{l}$
21:     Return $(x,\mathcal{P})$
Observe that these computations need to broadcast the global information of $u_{\min}(x,\mathcal{P})$, $x_{i_{\min}}(x,\mathcal{P})$ and $i_{\max}^{+}(x,\mathcal{P})$. For this purpose, we propose an algorithm based on message exchanges among neighbors. Agent $i$’s neighborhood $\mathcal{N}_{i}$ is defined as the set of agents whose partitions are adjacent to $\mathcal{P}_{i}$ as follows:
$\mathcal{N}_{i}=\{j\in\mathcal{N}\setminus\{i\}|{\rm adj}(\mathcal{P}_{i},\mathcal{P}_{j})\},$ (21)
where ${\rm adj}(\mathcal{P}_{i},\mathcal{P}_{j})=(\exists(c_{1},c_{2})\in\mathcal{E}|c_{1}\in\mathcal{P}_{i},c_{2}\in\mathcal{P}_{j})$ returns true if $\mathcal{P}_{i}$ and $\mathcal{P}_{j}$ are adjacent.
Algorithm 1 shows how to share global information, for example, $u_{\min}$, by constructing a communication tree in a distributed manner. The algorithm can be triggered by any agent $i$. The goal of the algorithm is to find agent $D_{i}$ that is the only parent of $i$ in the communication tree. Prior to executing the algorithm, each agent initializes its memorized utility $\hat{u}_{i}$ with its own utility $u_{i}$. Also, we set $\mathcal{\hat{N}}_{i}=\mathcal{N}_{i}$. Then firstly, agent $i$ communicates with its neighbors $\mathcal{\hat{N}}_{i}$ to find the agent $j_{{\rm lmin}}$ that has the locally minimum utility among them (Lines 2-5). If $i$ itself is $j_{{\rm lmin}}$, $i$ points to itself only when it has not pointed any other agents (Lines 7-9). If $j_{{\rm lmin}}$ is not either $i$ nor the current $D_{i}$, $i$ points to $j_{{\rm lmin}}$ (Lines 10-12). At the same time, $i$ also updates its utility in memory to serve as a proxy to the root ($i_{\min}$) in further communications (Line 13). Lastly, the algorithm of the neighborhood is triggered recursively until the convergence (Line 16). In the convergence, a tree rooted at $i_{\min}$ is formed through the chain of pointers $D_{i}$.
Figure 4 shows how the communication tree is constructed in the case of Example 1. In convergence, all the agents in the tree have a consensus on $i_{\min}$ (and $u_{\min}$), which is the root of the tree. The communication tree plays an important role not only in sharing global information but also in algorithm scalability. As we mentioned, the computation of Equation (13) is NP-hard and is not scalable in the size of the neighborhood $n_{i}$. To address this issue, we define a fixed size neighborhood, $\mathcal{N}^{2}_{i}=\{i,D_{i}\}$, where each agent $i$ has only its parent $D_{i}$ as a single neighbor and therefore $n_{i}=2$. We also update the definition of adjacency function to maintain consistency with $\mathcal{N}^{2}_{i}$ as ${\rm adj}(\mathcal{P}_{i},\mathcal{P}_{j})=(\exists(c_{1},c_{2})\in\mathcal{E}|c_{1}\in\mathcal{P}_{i},c_{2}\in\mathcal{P}_{j},j\in\mathcal{N}^{2}_{i})$. The distance function ||$\cdot$|| ⋅ | also needs to be redefined only on the edges in the communication tree, $\mathcal{E}\setminus\{(c_{1},c_{2})\in\mathcal{E}|c_{1}\in\mathcal{P}_{i},c_{2}\in\mathcal{P}_{j},\neg{\rm adj}(\mathcal{P}_{i},\mathcal{P}_{j})\}$.
@@
Given these notations, we propose a distributed algorithm returning an equilibrium solution $(x,\mathcal{P})$. The algorithm is summarized in Algorithm 2. At each iteration, an agent is activated. The agent communicates and coordinates with her neighborhood to approach a neighborhood optimum. Firstly, all agents initialize $\mathcal{P}_{i}$ with a Voronoi partition based on geodesic distances Haumann et al. (2011) (Line 2). As far as the state is not $Z_{4}$, each agent iterates the algorithm and shares the necessary information, $u_{\min}$, $x_{i_{\min}}$ and $i_{\max}^{+}$ based on a consensus algorithm via neighborhood communication, as in Algorithm 1 (Lines 3-4).
@@
Then the algorithm picks an agent $i$ (Line 5) in a distributed fashion. This is, for instance, done by sharing ’done/undone’ status and agent IDs, to pick up the ’undone’ agent with the smallest ID. Alternatively, agents could probabilistically move (probability $\alpha$) or not (probability $1-\alpha$). Agent $i$ computes its neighborhood $\mathcal{N}^{2}_{i}=\{i,D_{i}\}$ (Line 6). If the state is $Z_{1}$, it continues to pick another agent until $i_{\max}^{+}$ is in $\mathcal{N}^{2}_{i}$ to prioritize the process of $i_{\max}^{+}$ (Line 7-8). The activated agent $i$ then checks whether the neighborhood $\mathcal{P}_{i}$ could accommodate another agent (Line 10).
@@
  • Step a: If $\mathcal{P}_{n_{i}}$ cannot accommodate another agent, agent $i$ computes a neighborhood optimum for $\mathcal{N}^{2}_{i}$ and implements the allocation (Line 11).
@@
  • Step b: If $\mathcal{P}_{n_{i}}$ can accomodate another agent, it computes new locations $x^{\prime}$ for agents $\mathcal{N}^{2}_{i}$ with an additional agent $l$ (Line 15). Then, the algorithm makes a new partition for $n_{i}+1=3$ agents by splitting $\mathcal{P}_{n_{i}}$ based on the new locations $x^{\prime}$ (Line 16). To allocate the new partition, the algorithm finds agent $\tilde{i}_{\min}\not\in\mathcal{N}^{2}_{i}$ that is the closest to $i_{\min}$ in $\mathcal{N}_{\mathcal{N}^{2}_{i}}$, that is the neighbors of $\mathcal{N}^{2}_{i}$ (Line 17). Then the partition closest to $\tilde{i}_{\min}$ and adjacent to $\mathcal{P}_{\tilde{i}_{\min}}$ is allocated to $l$ first (Line 18). The other partitions are allocated to $\mathcal{N}^{2}_{i}$ by, for example, the optimal transport algorithm (Line 19). Lastly, $\mathcal{P}_{l}$ is merged to the partition of the agent that is the closest and adjacent to $l$ (Line 20).
@@
Figure 5. A pathological non-convex discrete example. Different from grid spaces, nodes are connected by edges. Left: Voronoi partition. Middle: A move of agent ’a’ changes partitions outside of the neighborhood. Right: Changes in $\mathcal{P}$ are confined inside the neighborhood.
In both cases, agents allocate the partition to avoid collisions after they redevide the neighborhood. Note that the algorithm updates the partition $\mathcal{P}_{j}$ of only neighbors $j\in\mathcal{N}^{2}_{i}$ and does not affect the other agents outside of $\mathcal{N}^{2}_{i}$ (Figure 5). This guarantees the convergence of the algorithm even in non-convex discrete settings as in Theorem 5.
The complete proof of Theorem 5 is in the supplementary material iwa (2024) due to the space limit. Briefly, it proves that the following potential function always increases for each iteration of the algorithm. Since the solution space is finite or compact, then the algorithm terminates.
@@
Also, the proof sketch of the approximation ratio is as follows. Because of the submodularity of $G$ and Equation (18), adding agents in optimal allocation $x^{*}$ to the same number of agents allocated by the algorithm does not make $G$ double. Formally, $G(x^{*})\leq G(x,x^{*})\leq 2G(x)$.
As shown in Theorem 5, the algorithm achieves the approximation ratio of 2, by communicating the minimum degree of information, $u_{\min}$, $i_{\max}^{+}$ and $x_{i_{\min}}$. The agent with the maximal gain from an additional agent in his region $i_{\max}^{+}$ is used to focus on the area to be reallocated in Line 7 of Algorithm 2. The minimum utility $u_{\min}$ is the key to proving the approximation ratio based on Equation (18). The location $x_{i_{\min}}$ is required to extend the algorithm for 1-dimensional setting in Marden (2016) to more general settings. Note that Algorithm 2 is not the only one that achieves the approximation ratio of 2. The required information also depends on the algorithm and $u_{\min}$, $i_{\max}^{+}$ and $x_{i_{\min}}$ are not the necessary conditions of the approximation ratio.
The algorithm works even better in a special case. Let $\mathcal{C}_{+}\subseteq C$ be a set of all important points such that $v_{c}=1,\leavevmode\nobreak\ \forall c\in\mathcal{C}_{+}$. Then other points are less important as $v_{c^{\prime}}\ll 1,\leavevmode\nobreak\ \forall c^{\prime}\in C\setminus\mathcal{C}_{+}$. The algorithm guarantees an optimal solution when agents can cover all the important points as follows.

### 08 Apr 2024 v3

@@ 3.1. Key Notions
Note that determining the locations $y_{1},\dots y_{k}$ maximizing Equation 13 is actually an NP-hard problem since it consists in general in solving a multiagent coverage problem.
We denote $\mathcal{P}_{ij}=\mathcal{P}_{i}\cup\mathcal{P}_{j}$. Informally, we say that an allocation is neighborhood optimal when no coalition of agents within the same neighborhood could reallocate themselves in a way which would improve the global objective, and would not benefit enough from accommodating a third agent in their combined regions. We will make these notions precise below. But before going into the details, we state our main result:
Theorem 4 (Convergence with Performance guarantee).
Under our communication model, Algorithm 1 terminates in a neighborhood optimum allocation. Its approximation ratio is 2.
Note that this approximation ratio of 2 is equal to the lower bound predicted by Equation 8 in case of full information ($f=n$). In our case it is achieved by only requiring agents to communicate beyond their neighborhood (i) the utility and location of the worst-off agent, and (ii) the identity of the agent which would contribute the most to the global objective by adding an agent in her region. In case of ties a deterministic choice mechanism is used, for instance based on agents’ id. Hence, the following notations will be useful:
$\left.\begin{array}[]{ll}i_{\max}^{+}(x,\mathcal{P})=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\mathop{\rm argmax\leavevmode\nobreak\ }\limits}\leavevmode\nobreak\ \!\!M_{1}(x_{i};\mathcal{P}_{i})&V(x,\mathcal{P})=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\max}\leavevmode\nobreak\ M_{1}(x_{i};\mathcal{P}_{i})\\ i_{\min}(x,\mathcal{P})=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\mathop{\rm argmin\leavevmode\nobreak\ }\limits}\leavevmode\nobreak\ \!\!u_{i}(x;\mathcal{P})&u_{\min}(x,\mathcal{P})=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\min}\leavevmode\nobreak\ u_{i}(x;\mathcal{P})\end{array}\right.$
@@
(Ex. 1, cont.). We have $i_{\min}=a$, since $u_{a}=1$. Furthermore, the agents which would contribute to $G(x)$ the most from adding a single agent within their regions are agents $e$ and $d$. For instance, adding an agent (depicted as ’+’) would induce $G$ of $6.5$ in her region, thus $M_{1}(x_{e},\mathcal{P}_{e})\simeq 6.5-5=1.5$ (See Figure 3 for the illustration). We assume $i^{+}_{\max}=d$ by lexicographic tie-breaking.
@@
Figure 3. Agent $e$ would contribute to $G(x)$ the most from the addition of an agent (depicted as ‘+’) in her region
3.2. High Level Description of the Algorithm
The principle of our algorithm follows Marden (2016), but is adapted to the 2D setting. Essentially, the idea is for neighboring agents to compare what they would gain by optimizing over their combined regions, or by optimizing over their combined regions with a third agent (which is necessarily at least as good). When the difference between these two situations is significant enough (larger than the utility of the worst-off agent), the decision is to make room for a third agent. This way, the space left open can eventually be filled.
As mentioned before, agents will communicate a limited amount of information. Spreading the information of $i_{\min}$ and $i^{+}_{\max}$ within the system can be achieved by standard distributed algorithms, eg. flooding and gossip protocols Wattenhofer (2016). At the end of this process, a communication spanning tree is built. Figure 4 shows a communication tree obtained for Example 1. In our algorithm, interaction takes place in the neighborhood defined by such communication trees, which are updated when required. In the following we denote by parent(i,$\mathcal{P}$) the parent of agent $i$, by neigh$(S,\mathcal{P})$ the neighbors of a set of agents $S$, and by ${E}(\mathcal{P})$ the set of all pairs of neighboring agents. All these notions are understood as restricted to the current communication tree (in its undirected version for the notion of neighborhood).
As in Marden (2016), it is useful to classify the solution $(x,\mathcal{P})$ into 4 states. First, the solution space is divided into the following two states $Z_{1}$ and $Z_{2}$, by checking if $i_{\max}^{+}$ would gain more from adding an agent in her region:
@@ 3.2. High Level Description of the Algorithm
In $Z_{2}$, no single agent would contribute enough from accommodating $i_{\min}$. Next the solutions in $Z_{2}$ are further classified depending on whether integrating a further agent in the neighborhood of two agents could induce a significant enough marginal gain. In the following state $Z_{3}$, it is not the case:
@@
$M_{3}(\emptyset,\mathcal{P}_{ij})-M_{2}(\emptyset,\mathcal{P}_{ij})\leq u_{\min}(x,\mathcal{P}),\leavevmode\nobreak\ \forall(i,j)\in{E}(\mathcal{P})\}$ (16)
Equation 16 means that the gain in the neighborhood optimum by accommodating $i_{\min}$ in $\mathcal{P}_{ij}$ cannot be larger than $u_{\min}$.
@@
$u_{i}(x,\mathcal{P})+u_{j}(x,\mathcal{P})=M_{2}(\emptyset,\mathcal{P}_{ij}),\leavevmode\nobreak\ \forall(i,j)\in{E}(\mathcal{P})\}$ (17)
@@
(Ex. 1, cont.) The allocation $x$ depicted is in state $Z_{1}$ since $V(x,\mathcal{P})=1.5>1=u_{a}$, with $i_{\min}=a$.
@@
Figure 4. (Left): the neighborhood graph in Example 1. The numbers represent utilities. (Right): a communication tree in convergence. In this case, agent ‘a’ is $i_{\min}$ and therefore the root of the tree.
@@
4:         Build communication tree
5:         Communicate $u_{\min}(x,\mathcal{P})$, $x_{i_{\min}}(x,\mathcal{P})$ and $i_{\max}^{+}(x,\mathcal{P})$
6:         if $(x,\mathcal{P})\in Z_{1}$ then
7:              $i\leftarrow i_{\max}^{+}$
8:         else
9:              pick $i\in\mathcal{N}$
10:         $j\leftarrow parent(i,\mathcal{P})$
11:         if $i_{\min}(x,\mathcal{P})=j$ or
12:            $M_{3}(\emptyset,\mathcal{P}_{ij})-M_{2}(\emptyset,\mathcal{P}_{ij})\leq u_{\min}(x,\mathcal{P})$ then
13:               $x_{\{i,j\}}\leftarrow B_{2}(\emptyset,\mathcal{P}_{ij})$ (Step a)
14:              $\mathcal{P}_{k}\leftarrow\mathcal{V}_{k}(x^{\prime},\mathcal{P}_{ij}),\leavevmode\nobreak\ \forall k\in\{i,j\}$
15:         else
16:              Consider a virtual agent $l$ (Step b)
17:              $x^{\prime}\leftarrow B_{3}(\emptyset,\mathcal{P}_{ij})$
18:              $\mathcal{P}^{\prime}_{k}\leftarrow\mathcal{V}_{k}(x^{\prime},\mathcal{P}_{ij}),\leavevmode\nobreak\ \forall k\in\mathcal{N}^{\prime}_{i}=\{1,2,3\}$
@@
19:              $\tilde{i}_{\min}=\underset{\begin{subarray}{c}k\in neigh(\{i,j\},\mathcal{P})\end{subarray}}{\mathop{\rm argmin\leavevmode\nobreak\ }\limits}|x_{j}-x_{i_{\min}}|$
@@
20:              $k_{l}=\underset{\begin{subarray}{c}k\in neigh(\tilde{i}_{\min}\end{subarray},\mathcal{P})\cap\mathcal{N}^{\prime}_{i}}{\mathop{\rm argmin\leavevmode\nobreak\ }\limits}|x^{\prime}_{k}-x_{\tilde{i}_{\min}}|,\leavevmode\nobreak\ x_{l}\leftarrow x^{\prime}_{k_{l}}$
21:              $x_{\{i,j\}}\leftarrow x^{\prime}\setminus\{x_{l}\}$
@@
22:              $i_{+}\leftarrow\underset{\begin{subarray}{c}k\in neigh(l,\mathcal{P})\cap\{i,j\}\end{subarray}}{\mathop{\rm argmin\leavevmode\nobreak\ }\limits}\leavevmode\nobreak\ |x_{j}-x_{l}|,\leavevmode\nobreak\ \mathcal{P}_{i_{+}}\leftarrow\mathcal{P}_{i_{+}}\cup\mathcal{P}_{l}$
23:     Return $(x,\mathcal{P})$
@@
Given this classification, we propose a distributed algorithm returning an equilibrium solution $(x,\mathcal{P})$. The algorithm is summarized in Algorithm 1. Firstly, all agents initialize $\mathcal{P}_{i}$ with a Voronoi partition based on geodesic distances Haumann et al. (2011) (Line 2). As far as the state is not $Z_{4}$, agents build a communication tree and share the necessary information, $u_{\min}$, $x_{i_{\min}}$ and $i_{\max}^{+}$ based on a consensus algorithm via neighborhood communication. (Lines 3-5).
@@
If the algorithm is in state $Z_{1}$, the agent $i^{+}_{\max}$ is picked (Line 9), otherwise a random agent is picked. Note that in a distributed fashion, this could be done for instance by sharing ‘done/undone’ status and agent IDs, to pick up the ‘undone’ agent with the smallest ID. Alternatively, agents could probabilistically move (probability $\alpha$) or not (probability $1-\alpha$). Agent $i$ retrieves her parent (denoted $j$) in the current communication tree. The activated agent $i$ then checks whether the combined regions could accommodate another agent (Line 12).
@@
  • Step a: If $\mathcal{P}_{ij}$ cannot accommodate another agent or $i_{\min}$ belongs to this neighborhood, agent $i$ computes a neighborhood optimum for the pair of agents $(i,j)$ and the allocation is implemented (Line 13).
@@
  • Step b: If $\mathcal{P}_{ij}$ can accomodate another agent, it computes new locations $x^{\prime}$ for agents $i$ and $j$, together with an additional agent $l$ (Line 17). Then, the algorithm computes a new partition for these $3$ agents by splitting $\mathcal{P}_{ij}$ based on the new locations $x^{\prime}$ (Line 18). To allocate the new partition, the algorithm finds the agent $\tilde{i}_{\min}\not\in\{i,j\}$ that is the closest to $i_{\min}$ in the communication tree. Then the partition closest to $\tilde{i}_{\min}$ and adjacent to $\mathcal{P}_{\tilde{i}_{\min}}$ is allocated to $l$ first (Line 20). The other partitions are allocated to $i$ and $j$, for example by an optimal matching algorithm so as to minimize the moves of the agents (Line 21 Crouse (2016)). Lastly, $\mathcal{P}_{l}$ is merged to the partition of the agent $i_{+}$ that is the closest and adjacent to $l$ (Line 22).
@@
Figure 5. A pathological non-convex discrete example. Different from grid spaces, nodes are connected by edges. Left: Voronoi partition. Middle: A move of agent $a$ changes partitions outside of the neighborhood. Right: Changes in $\mathcal{P}$ are confined inside the neighborhood.
In both cases, agents allocate the partition to avoid collisions after they redevide the neighborhood. Note that the algorithm updates the partition $\mathcal{P}_{j}$ of only neighbors $(i,j)\in E$ and does not affect the other agents outside of $(i,j)$ (Figure 5). This avoids issues with the convergence of the algorithm in non-convex discrete settings, as discussed in Yun and Rus (2014).
[…]
Figure 6. After step b, $c$ and $d$ have reallocated themselves in $\mathcal{P}_{c}\cup\mathcal{P}_{d}$ assuming a third agent would join, leaving the empty position (unfilled) as close as possible from $a$ ($i_{\min}$)
Let us illustrate this step on our example (Figure 6). For instance, assuming $d$ is picked as the $i^{+}_{\max}$ agent, this agent interacts with $c$, her parent in the communication graph. Upon evaluating the optimal partition with a third agent, they assess that the gain is higher than $u_{\min}$. As a result, they move and make room for a third agent. As discussed before, note that the region is not updated with respect to that of agent $b$ at this stage.
Due to the space limitation, the complete proof of Theorem 4 is in the supplementary material Iwase et al. (2024). It follows the line of reasoning described in Marden (2016) but adapted to our setting. Briefly, it proves that the following potential function always increases for each iteration of the algorithm. Since the solution space is finite or compact, then the algorithm terminates.
@@
Also, the proof sketch of the approximation ratio is as follows. Because of the submodularity of $G$ and Equation  16, adding agents in optimal allocation $x^{*}$ to the same number of agents allocated by the algorithm does not make $G$ double. Formally, $G(x^{*})\leq G(x,x^{*})\leq 2G(x)$.
3.3. Special Case
It is interesting to observe that the algorithm provides an optimal solution in the special case where there are exactly as many points of interest as agents. Let $\mathcal{C}_{+}\subseteq C$ be a set of all important points such that $v_{c}=1,\leavevmode\nobreak\ \forall c\in\mathcal{C}_{+}$. Then other points are less important as $v_{c^{\prime}}\ll 1,\leavevmode\nobreak\ \forall c^{\prime}\in C\setminus\mathcal{C}_{+}$. The algorithm guarantees an optimal solution when agents can cover all the important points as follows.